OpenStudy (anonymous):

Can anybody please help me solve this..i already have the answer..i just need to know the steps.. Find the solution to the equation 64^(x – 4) = 4^(2x).

OpenStudy (anonymous):

x=12..my question is why does it equals to 12...what are the steps taken to get the answer

OpenStudy (anonymous):

match the bases then solve for x

OpenStudy (anonymous):

how can you rewrite 64 ?

OpenStudy (anonymous):

what do you mean by match the bases? 4^16 = 64 ..how will we implement this?

OpenStudy (anonymous):

4^3(x-4)=4^(2x)

OpenStudy (anonymous):

4^3(x-4)=4^(2x) 3(x-4)= 2x 3x-12 = 2x -12= -x -x/-1= -12/-1???

OpenStudy (anonymous):

therefore x=12?

OpenStudy (anonymous):

3(x-4) = 2x x-(2/3)x = 4 (1/3)x = 4 so x= 12

OpenStudy (anonymous):

yep.. its easier when the bases are the same

OpenStudy (anonymous):

hmm..i see..

OpenStudy (anonymous):

thanks alot for the help...

OpenStudy (anonymous):

if not.. then you would have to go logging

OpenStudy (anonymous):

oooh ok..thats what i was thinking..that i have to do logging on this one too..and just the though was very intimidating..lol

OpenStudy (anonymous):

thought*

OpenStudy (anonymous):

its not that bad.. just practice

OpenStudy (anonymous):